# Wave Function Collapse, Lorentz Invariance, and the Third Postulate of Relativity

### v2

@@ Abstract
In this work it is shown that this modified Schrödinger equation is Lorentz invariant. It is further argued that the additional spacetime structure that it requires provides a way to implement the assumption that spacelike-separated operators (and measurements) commute, and that this assumption of local commutativity should be regarded as a third postulate of relativity.
@@ I Introduction
Gisin’s work was one of a number of proposed stochastic modifications of the Schrödinger equation aimed at resolving the measurement problem[2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Most of these proposals are designed to collapse the state vector to either an approximate position state or to an energy eigenstate. These attempted solutions have often been met with skepticism because they introduce new, ad hoc physical constants and imply small violations of conservation laws. A recent work has shown how to eliminate these problematic features[16]. It is based on the idea that wave function collapse is induced by the elementary interactions that establish correlations between physical systems, and was motivated by the fact that these correlating interactions play a central role both in the measurement of physical quantities and in the instantiation and transmission of physical information. In this work it will also be shown that this proposed equation maintains Lorentz invariance.
Since the proposed equation is Lorentz invariant it is reasonable to ask whether it can be considered “genuinely relativistic”. In order to answer this question it is necessary to deal with the nonlocal nature of wave function collapse.
@@
In the conventional view relativistic spacetime is a four-dimensional manifold with a Lorentzian metric that defines a light cone structure. This framework implements Einstein’s two postulates for relativity by prohibiting the assignment of an absolute temporal order to events that are separated by a spacelike interval. Why, then, did the advent of quantum theory necessitate the introduction of a third postulate to maintain this prohibition? Prior to the development of quantum theory the relativistic prohibition on temporal order was strongly associated with the presumption that no physical processes could propagate outside the light cone. But the nonlocal correlations implied by quantum theory strongly suggest that there are physical effects that propagate across spacelike intervals. To “explain” why these effects do not result in any manifest inconsistencies with relativity the new theory simply ruled by fiat that they had to respect the prohibition on temporal ordering. As noted, this was done by requiring that spacelike-separated operators commute. For the reasons mentioned a few paragraphs back there was no inclination to consider the possibility that spacetime possessed additional structure. In fact, the assumption of local commutativity was often conflated with the limitation on the speed of light. The following quote from Gell-Mann, Goldberger, and Thirring illustrates this point[23]:
@@
As suggested by Weinberg’s very careful characterization of local commutativity described above, the kind of conflation demonstrated in the quotation is simply wrong. The assumption that spacelike-separated operators commute is made to insure that any effects on physical systems that do propagate across spacelike intervals do not transmit any physical information. It involves the (at least) implicit recognition that there are nonlocal effects, and that there is a need to regulate them. But, this kind of regulation ought to be explained, and not simply imposed by fiat.
As mentioned above Einstein’s postulate about the invariance of the speed of light is implemented by attributing a light cone structure to spacetime. In other words, this postulate is explained as a consequence of the fundamental nature of spacetime. Should we not then also consider adding structure to spacetime to explain how nonlocal effects propagate and how they can be regulated to maintain Lorentz invariance? This is exactly what stochastic collapse equations do. By assuming a foliation of spacetime into spacelike surfaces, and invoking a stochastic process (or processes) they provide the desired explanation. By making wave function collapse and the Born rule follow from the fundamental equation of the theory, they provide a coherent logical structure for contemporary physics.
To assess the extent to which these proposals can be judged to be genuinely relativistic we can start by stipulating that the primary criterion for a relativistic theory is that the laws of physics have the same form in all inertial frames. In particular, this means that they must be Lorentz invariant. Assertions that a proposed account can be judged as “genuinely relativistic” only if it does not add structure beyond the standard relativistic metric cannot be accepted as a priori truths. They face a burden of proof that must take into account the nonlocal aspects of quantum theory. In order to reconcile relativity and quantum theory one must consider the possibility that both these pillars of contemporary physics will have to be modified. The evaluation of any particular proposal will then turn on whether it is Lorentz invariant and whether the Lorentz invariance follows from fundamental physical principles in a natural manner. It will be argued here that the proposal in [16] meets these criteria.
The next section describes the basic assumptions employed by nonlinear, norm-preserving stochastic collapse equations and illustrates how they work. Section III reviews the proposal of [16], which eliminates the need to introduce new ad hoc physical constants and insures that conservation laws are respected in individual experiments. It then goes on to demonstrate that it meets a critical test of Lorentz invariance in a very natural way. Section IV addresses the question of what it means for a theory to be ”genuinely relativistic.”
@@ II Stochastic Collapse Equations
This section first provides an explanation of why stochastic collapse equations need to add structures to spacetime, specifically a preferred foliation associated with a stochastic process (or processes). This is followed by an illustration of how these dynamic equations generate collapse in accord with the Born rule. The literature cited earlier contains more general and formal demonstrations of how these equations work; a proof that the equation described in Section III entails collapse in conformity to the Born rule has been presented in [16]. The purpose here is just to provide an intuitive understanding based on a simplified case that captures the essential features. The idea is to show that it is not necessary to paste the measurement postulates onto quantum theory in an ad hoc manner, but rather that they follow in a very natural way from a relatively simple modification of the Schrödinger equation that takes into account the fundamentally probabilistic nature of quantum theory as we currently understand it.
@@ II.2 Nonlinear, Norm-Preserving Stochastic Collapse Equations
With this background we can now review the proposed equation described in [16], and see how it maintains Lorentz invariance.
@@ III Interaction-Induced Wave Function Collapse
To establish that the expression for the denominator in $\gamma_{jk}$, $||\langle\,\psi|\mathbf{{V}_{jk}}|\,\psi\rangle||_{max}$, is well defined note that the collapse equation can be integrated both backward and forward in time. Because the Hamiltonian is independent of time it is clear that the associated Scrödinger equation is integrable. The fact that the stochastic operator is unbiased implies that averaging over all of the unravellings of the collapse equation will reproduce the value obtained by the Scrödinger equation integral. Observe also that the stochastic changes in the denominator are matched by those in the numerator; so their overall effect on $\gamma_{jk}$ is to leave it essentially unchanged. It is, of course, not necessary to actually carry out the integration since $\gamma_{jk}$ is designed to integrate to a value of order, $1$, over the course of the interaction. With this understanding the variable rate parameter is defined as follows:
@@
The period over which $\gamma_{jk}$ is significantly different from zero depends on the initial conditions of the interaction. However, the time during which the vast majority of the momentum and energy are exchanged (and the amplitude is transferred) depends essentially on the maximum interaction energy:
@@
In other words, the common textbook example of a particle in a box ignores the small amount of entanglement that results from the interaction between the particle, the box and whatever apparatus was used to prepare the system. The interactions involved in both the preparation and measurement of quantum systems are conservative interactions. When these interactions induce branching of the wave function they lead to a different distribution of conserved quantities among the interacting systems in the various branches, but they do not alter the total amount (up to normalization of the branches.) Based on these kinds of observations a number of articles in recent years have demonstrated that conservation laws hold exactly in individual instances, and not just when applied to ensembles of identical measurement situations. In addition to [16] these include [39, 30, 31, 32, 33, 34, 35].
In discussing the status of conservation laws with reference to stochastic collapse equations there are two issues that must be addressed. First, some of the collapse proposals imply small violations of conservation laws even with regard to statistical averages. For example, consider the most widely known collapse proposal, the continuous spontaneous localization (CSL) model[36, 37]. It implies small violations of energy conservation as described in [38]. One of the main reasons for these violations is that this proposal uses a stochastic field, $\xi(x,t)$, which is a function of both position and time, as opposed to a single, global stochastic process, $\xi(t)$, which is a function only of time. The stochastic field induces spatial variations in the wave function - in effect, pumping energy into it. Other proposals, such as one of the variants discussed in [1] use a single, global stochastic process, and base the stochastic operator on the Hamiltonian. In this way it maintains energy conservation on average, but still violates it in individual instances.
@@
This review of the proposal in [16] is intended as background for the demonstration that 10 is Lorentz invariant. The general question of Lorentz invariance hangs over any proposal that deals with the measurement problem. But there is a particular concern that arises in connection with the type of collapse equation discussed here because they are formulated in a preferred reference frame. There is essentially zero probability that any reference frame in which one chooses to analyze a collapse process will coincide with the preferred frame. The critical question is whether the magnitude of the collapse effects is independent of the frame in which they are considered.
@@
IV Reconciling Relativity and Quantum Nonlocality
One common attitude is that no proposal that requires adding structure to spacetime beyond a Lorentzian metric can be considered to be genuinely relativistic. It implies that any new physics must work within the limited framework that was erected in a classical context prior to the development of quantum theory.
To evaluate this viewpoint it is necessary to take into account the fundamentally probabilistic and nonlocal nature of quantum theory as we currently understand it. As argued in Section I these characteristics necessitated the introduction of an additional postulate for theoretical physics - local commutativity. Because the new theory was such a large break from the past, and because many of its consequences only became apparent over time the central role that this additional postulate plays in contemporary physics has not been fully appreciated.
The Lorentzian metric and light cone structure that we use to characterize relativistic spacetime were adopted in order to provide a fundamental physical explanation for Einstein’s two original postulates for relativity. The assumption of local commutativity deals with the novel aspects of quantum theory by forbidding the transmission of stable physical information across spacelike intervals. It is intended to prevent any manifest violations of relativity and should be regarded as a third postulate. But, just as the earlier structures were designed to explain how the original postulates are enforced we should be willing to entertain the possibility of adding structure to explain how this third postulate is enforced. That is exactly what a spacetime foliation associated with a stochastic process does.
@@ IV Reconciling Relativity and Quantum Nonlocality
The adoption of this sort of additional stochastic structure transfers the probabilistic nature of quantum theory from the macroscopic to the elementary level. This turns what were ad hoc postulates at the macro level into straightforward consequences of the mathematical description of elementary processes. The probabilistic nature of elementary correlating interactions also explains why a preferred reference frame (or spacetime foliation) remains hidden, and explains why the description of spacetime at the macroscopic level is limited to the standard relativistic account involving just a Lorentzian metric. This is why accounts involving these sorts of additional stochastic structures ought to be regarded as genuinely relativistic.
@@
Because the proposal described here is formulated in a nonrelativistic framework it must be considered as incomplete. But since it can, arguably, be regarded as Lorentz invariant there do not appear to be any serious conceptual obstacles to a full relativistic account. Such an account would encompass quantum field theory, and might well offer solutions to some of the key problems with this more limited proposal. As argued in [16] there is reason to believe that such an extension could resolve the small discrepancies with energy conservation that are implied by equation 10. One might also hope that a relativistic account might provide a solution to the tails problem (mentioned in a footnote in section II). With the transition from distinguishable to indistinguishable particles the suppression of the very small amplitudes in the wave function tail below the level of vacuum fluctuations might point the way to a resolution. But these are just speculations that must be addressed in future work.
@@
Acknowledgements: I would like to thank Nicolas Gisin and Flavio Del Santo for useful discussions.

### v3

@@ Abstract
In this work it is shown that the stochastic modification to the Schrödinger equation is Lorentz invariant. It is further argued that the additional spacetime structure that it requires provides a way to implement the assumption that spacelike-separated operators (and measurements) commute, and that this assumption of local commutativity should be regarded as a third postulate of relativity.
@@ I Introduction
Gisin’s work was one of a number of proposed stochastic modifications of the Schrödinger equation aimed at resolving the measurement problem[2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Most of these proposals are designed to collapse the state vector to either an approximate position state or to an energy eigenstate. These attempted solutions have often been met with skepticism because they introduce new, ad hoc physical constants and imply small violations of conservation laws. A recent work has shown how to eliminate these problematic features[16]. It is based on the idea that wave function collapse is induced by the elementary interactions that establish correlations between physical systems, and was motivated by the fact that these correlating interactions play a central role both in the measurement of physical quantities and in the instantiation and transmission of physical information. In this work it will be shown that the proposed modification to the equation is Lorentz invariant. [^1] Of course, the nonrelativistic Schrödiger equation is not, itself, invariant. What will be shown here is that the stochastic modification to the equation is Lorentz invariant.
The fact that wave function collapse can be described in a Lorentz invariant manner calls into question our current understanding of relativity. What do these nonlocal, but Lorentz invariant effects imply about the ontology of spacetime?
@@
In the conventional view relativistic spacetime is a four-dimensional manifold with a Lorentzian metric that defines a light cone structure. This framework implements Einstein’s two postulates for relativity by prohibiting the assignment of an absolute temporal order to events that are separated by a spacelike interval. Why, then, did the advent of quantum theory necessitate the introduction of a third postulate to maintain this prohibition? Prior to the development of quantum theory the relativistic prohibition on temporal order was typically associated with the presumption that no physical processes could propagate outside the light cone. But the nonlocal correlations implied by quantum theory strongly suggest that there are physical effects that propagate across spacelike intervals. To “explain” why these effects do not result in any manifest inconsistencies with relativity the new theory simply ruled by fiat that they had to respect the prohibition on temporal ordering. As noted, this was done by requiring that spacelike-separated operators commute. For the reasons mentioned a few paragraphs back there was no inclination to consider the possibility that spacetime possessed additional structure. In fact, the assumption of local commutativity was often conflated with the limitation on the speed of light. The following quote from Gell-Mann, Goldberger, and Thirring illustrates this point[23]:
@@
As suggested by Weinberg’s very careful characterization of local commutativity described above, the kind of conflation demonstrated in the quotation is simply wrong. The assumption that spacelike-separated operators commute is made to insure that any effects on physical systems that do propagate across spacelike intervals do not transmit any physical information. It involves the (at least) implicit recognition that there are nonlocal effects, and that there is a need to regulate them. If this were not the case there would be no need to make an additional assumption. But, this kind of regulation ought to be explained, and not simply imposed by fiat.
As mentioned above Einstein’s postulate about the invariance of the speed of light is implemented by attributing a light cone structure to spacetime. In other words, this postulate is explained as a consequence of the fundamental nature of spacetime. Should we not then also consider adding structure to spacetime to explain how nonlocal effects propagate and how they can be regulated to maintain Lorentz invariance? This is exactly what stochastic collapse equations do. By assuming a foliation of spacetime into spacelike surfaces, and invoking a stochastic process (or processes) they provide the desired explanation. By making wave function collapse and the Born rule follow from the fundamental equation of the theory, they provide a more coherent logical structure for contemporary physics.
In order to provide a fully coherent logical framework for contemporary theory it is crucial that any proposed modifications to the fundamental equations be Lorentz invariant. The goals of this work are to show that the proposal of [16] meets this criterion, and to examine its implications for our understanding of relativity. The next section describes the basic assumptions employed by nonlinear, norm-preserving stochastic collapse equations and illustrates how they work. Section III reviews the proposed modification, showing how it eliminates the need to introduce new ad hoc physical constants and insures that conservation laws are respected in individual experiments. It then goes on to demonstrate that it meets a critical test of Lorentz invariance in a very natural way. Based on the demonstration of Lorentz invariance Section IV examines the implications for our understanding of relativity and spacetime ontology.
@@ II Stochastic Collapse Equations
This section first provides an explanation of why stochastic collapse equations need to add structure to spacetime, specifically a preferred foliation associated with a stochastic process (or processes). This is followed by an illustration of how these dynamic equations generate collapse in accord with the Born rule. The literature cited earlier contains more general and formal demonstrations of how these equations work; a proof that the equation described in Section III entails collapse in conformity to the Born rule has been presented in [16]. The purpose here is just to provide an intuitive understanding based on a simplified case that captures the essential features. The idea is to show that it is not necessary to paste the measurement postulates onto quantum theory in an ad hoc manner, but rather that they follow in a very natural way from a relatively simple modification of the Schrödinger equation that takes into account the fundamentally probabilistic nature of quantum theory as we currently understand it.
@@ II.2 Nonlinear, Norm-Preserving Stochastic Collapse Equations
With this background we can now review the proposed equation described in [16], and see how the stochastic modification maintains Lorentz invariance.
@@ III Interaction-Induced Wave Function Collapse
To establish that the expression for the denominator in $\gamma_{jk}$, $||\langle\,\psi|\mathbf{{V}_{jk}}|\,\psi\rangle||_{max}$, is well defined note that the collapse equation can be integrated both backward and forward in time. Because the Hamiltonian is independent of time it is clear that the associated Schrödinger equation is integrable. The fact that the stochastic operator is unbiased implies that averaging over all of the unravellings of the collapse equation will reproduce the value obtained by the Schrödinger equation integral. Observe also that the stochastic changes in the denominator are matched by those in the numerator; so their overall effect on $\gamma_{jk}$ is to leave it essentially unchanged. It is, of course, not necessary to actually carry out the integration since $\gamma_{jk}$ is designed to integrate to a value of order, $1$, over the course of the interaction. With this understanding the variable rate parameter is defined as follows:
@@
The period during which $\gamma_{jk}$ is significantly different from zero depends on the initial conditions of the interaction. However, the time during which the vast majority of the momentum and energy are exchanged (and the amplitude is transferred) depends essentially on the maximum interaction energy:
@@
In other words, the common textbook example of a particle in a box ignores the small amount of entanglement that results from the interaction between the particle, the box and whatever apparatus was used to prepare the system. The interactions involved in both the preparation and measurement of quantum systems are conservative interactions. When these interactions induce branching of the wave function they lead to a different distribution of conserved quantities among the interacting systems in the various branches, but they do not alter the total amount (up to normalization of the branches). Based on these kinds of observations a number of articles in recent years have demonstrated that conservation laws hold exactly in individual instances, and not just when applied to ensembles of identical measurement situations. In addition to [16] these include [35, 29, 30, 31, 32, 33, 34].
In discussing the status of conservation laws with reference to stochastic collapse equations there are two issues that must be addressed. First, some of the collapse proposals imply small violations of conservation laws even with regard to statistical averages. For example, consider the most widely known collapse proposal, the continuous spontaneous localization (CSL) model[36, 37]. It implies small violations of energy conservation as described in [38]. One of the main reasons for these violations is that this proposal uses a stochastic field, $\xi(x,t)$, which is a function of both position and time, as opposed to a single, global stochastic process, $\xi(t)$, which is a function only of time. The stochastic field induces spatial variations in the wave function - in effect, pumping energy into it. Other proposals, such as one of the variants discussed in [1] use a single, global stochastic process, and bases the stochastic operator on the Hamiltonian. In this way it maintains energy conservation on average, but still violates it in individual instances.
@@
This review of the proposal in [16] is intended as background for the demonstration that the action of the stochastic operator in 10 is Lorentz invariant. The general question of Lorentz invariance hangs over any proposal that deals with the measurement problem. But there is a particular concern that arises in connection with collapse equations of the type discussed here because they are formulated in a preferred reference frame. There is essentially zero probability that any reference frame in which one chooses to analyze a collapse process will coincide with the preferred frame. The critical question is whether the magnitude of the collapse effects is independent of the frame in which they are considered.
@@
IV Relativity, Nonlocal Quantum Effects, and Spacetime Ontology
The principle of relativity states that the laws of physics have the same form in all reference frames. Since the speed of light has been observed to be independent of the reference frame, a satisfactory theory should explain this invariance in a natural way without positing any unnecessary features.
Early works by Fitzgerald, Lorentz, Poincaré, and Larmor that were aimed at explaining the observed invariance assumed that electromagnetic waves propagated through a stationary aether. These researchers accounted for the experimental results by proposing that their measurement instruments - rods, clocks, and all matter - interacted according to the laws of electromagnetism which had been shown to be Lorentz invariant. The resulting behavior of the instruments made it impossible to pick out any special reference frame. Thus, Lorentz invariance became the principal defining characteristic of a relativistic theory. In his review of these early works Bell[39] illustrates how they established that
“if physical laws are Lorentz invariant…moving observers will be unable to detect their motion.”
The term, “motion”, of course, refers to motion with respect to the aether. What Einstein was able to show was that the hypothesis of an undetectable aether was superfluous; it added nothing to the explanation of the observed phenomena.
In the previous section it was shown that the stochastic modification to the Schrödinger equation proposed in [16] is Lorentz invariant, and that, therefore, the preferred frame that it assumes remains undetectable. What then distinguishes this situation from the pre-Einstein versions of relativity? The critical difference is that in this case the hidden spacetime feature plays a critical explanatory role. It provides a framework within which the basic dynamic equation of the theory can explain the nonlocal correlations described by Bell. Moreover, the fundamentally probabilistic nature of the collapse equation also explains why the preferred frame remains hidden.
What does this imply about our understanding of relativity? Following Einstein’s 1905 papers [40, 41] Minkowski presented his elegant mathematical description[42] which provided a very compelling account of spacetime structure. The Lorentzian metric defined a framework in which it appeared that all physical processes were constrained to propagate only within light cones. This provided an explanation for Einstein’s postulates in terms of the fundamental features of spacetime. The fact that it dovetailed beautifully with our intuitive notion that causal processes propagate through space in a continuous manner was very satisfying. When Einstein was able to generalize this picture to account for gravitational effects without invoking action-at-a-distance it cemented the idea that we had a complete account of spacetime ontology.
The advent of quantum theory radically altered this situation. Einstein was the first to recognize this, but it was Bell’s demonstration of the reality of nonlocal effects that has really forced us to reconsider the ontology of spacetime. It is time to acknowledge that the assumption of local commutativity constitutes a third postulate of relativity, and that, like Einstein’s two original postulates it should be explained in terms of spacetime structure. The nonlocal correlations should, of course, be explained in a natural way. The stochastic collapse equation described in Section III shows how this can be done.
@@ IV Relativity, Nonlocal Quantum Effects, and Spacetime Ontology
The adoption of the additional stochastic structure described in Section III transfers the probabilistic nature of quantum theory from the macroscopic to the elementary level. This turns what were ad hoc postulates at the macro level into straightforward consequences of the mathematical description of elementary processes. The probabilistic nature of elementary correlating interactions also explains why a preferred reference frame (or spacetime foliation) remains hidden, and explains why the description of spacetime at the macroscopic level is limited to the standard relativistic account involving just a Lorentzian metric.
It is important to note that although the additional spacetime structure posited by stochastic collapse equations remains hidden there is, nevertheless, experimental evidence for its existence. The numerous experiments confirming the nonlocal correlations identified by Bell strongly argue for some kind of connection across spacelike intervals. Additional confirmation is possible. The proposals offered to explain the correlations usually make some testable predictions. For example, the proposal in [16] predicts specific quantitative deviations from the linearity of standard quantum theory. These deviations are, of course, quite small. They are proportional to the square of the amplitude shifts, that is the ratio of interaction energy to total relativistic energy of the interacting systems. For two electrons separated by the Bohr radius these deviations from linearity would be about $10^{-9}$. Relevant experiments would be quite challenging, but they are not impossible.
@@
Because the proposal described here is formulated in a nonrelativistic framework it must be considered as incomplete. But since the proposed modification is Lorentz invariant there do not appear to be any serious conceptual obstacles to a full relativistic account. Such an account would encompass quantum field theory, and might well offer solutions to some of the key problems with this more limited proposal. As argued in [16] there is reason to believe that such an extension could resolve the small discrepancies with energy conservation that are implied by equation 10. One might also hope that a relativistic account might provide a solution to the tails problem (mentioned in a footnote in section II). With the transition from distinguishable to indistinguishable particles the suppression of the very small amplitudes in the wave function tail below the level of vacuum fluctuations might point the way to a resolution. But these are just speculations that must be addressed in future work.
@@
Acknowledgements: I would like to thank Nicolas Gisin and Flavio Del Santo for helpful discussions.